# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v2

@@ 5. From Classic Isoperimetry to Log-Sobolev
Assume that (1.5) holds. Then, for any $\alpha>0$,
@@
We first prove the first inequality in (5.1). Let $A$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ for $\nu=\nu_{X}=\nu_{Y}$ (given in (3.2)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [42, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
@@
$\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A^{\sqrt{n\tau}})/\nu_{n}(A)]}{n\sqrt{\tau}}$
$\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
@@
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 8 again, it holds that
$\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$
Hence, for any $\delta>0$,
$\bar{\Lambda}(\psi(\alpha,\tau)-\delta)$ $\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, it holds that $\psi(\alpha,\tau)>\alpha-\delta.$ for all sufficiently small $\tau$, which implies
$\bar{\Lambda}(\alpha-\delta)$ $\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
Letting $\tau\downarrow 0$ yields that
@@
$D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})$
$=\mathbb{E}_{\pi}[\alpha_{W}-\beta_{W}]$
$\leq\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]\mathbb{E}_{\pi}[\gamma_{W}^{2}]}.$
@@
We now need the HWI inequality for the Riemannian manifold setting [47, Corollary 20.13]. The original version for Euclidean spaces is due to Otto and Villani [38].
Theorem 17 (HWI Inequality [47, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that (1.5) holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
@@
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality; see [4, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $P_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\pi_{XW}:\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\,D(\pi_{X|W}\|\nu|\pi_{W})=\alpha}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}$
$=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (5.2) yields that
$\bar{\Lambda}(\alpha-\delta)$ $\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
That is, for $\alpha,\delta>0$,
$\bar{\Lambda}(\alpha)$ $\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$
Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields
$\bar{\Lambda}(\alpha)$ $\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Let $k$ be the dimension of $\mathcal{M}$. Then, the coarea formula (see e.g., [14, Exercise III.12]) states:
@@
$=\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}^{+}\{f^{\otimes n}\geq s\}{\text{\rm d}}s$
$\geq\eta_{n}\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s,$
@@
$\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s$
$=\int_{0}^{\infty}\int 1\{f^{\otimes n}\geq s,\,s\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}{\text{\rm d}}s$
$=\int(f^{\otimes n}-e^{n(D-\delta)})\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}$
$=\gamma_{1,n}-\gamma_{2,n},$
@@
$\gamma_{1,n}$ $:=\int f^{\otimes n}\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}$
$=\mu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\},$
@@
By the law of large numbers,
$\gamma_{1,n}$ $\to 1,\textrm{ as }n\to\infty,$
By the large deviations theory (or the asymptotic equipartition property in information theory),
@@
where
$D_{1-t}(\mu\|\nu)=-\frac{1}{t}\ln\int f^{1-t}d\nu,\;f={\text{\rm d}}\mu/{\text{\rm d}}\nu$
is the (1t)(1-t)-Rényi divergence of $\mu$ w.r.t. $\nu$.
Therefore,
$\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\to 1,\textrm{ as }n\to\infty,$
which implies
@@
(5.5) $\eta_{n}$ $\geq\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{I_{n}(\nu_{n}\{f^{\otimes n}\geq s\})}{\nu_{n}\{f^{\otimes n}\geq s\}}$
(5.6) $=\frac{I_{n}(a_{n})}{a_{n}},$
where $a_{n}:=\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}$, and (5.6) follows since $I_{n}(a)/a$ is non-increasing in $a$. Hence,
@@
$\int|\nabla f^{\otimes n}|\,{\text{\rm d}}\nu_{n}$ $\leq\sqrt{\int f^{\otimes n}\,{\text{\rm d}}\nu_{n}\,\cdot\int\frac{|\nabla f^{\otimes n}|^{2}}{f^{\otimes n}}\,{\text{\rm d}}\nu_{n}}$
(5.8) $=\sqrt{n\int\frac{|\nabla f|^{2}}{f}\,{\text{\rm d}}\nu}=\sqrt{nI(\mu\|\nu)}.$
Combining (5.4), (5.7), and (5.8) yields that
$\sqrt{I(\mu\|\nu)}\geq\bar{\Lambda}(D(\mu\|\nu)-\delta),\,\forall\mu.$
Noting that $\bar{\Lambda}$ is non-decreasing, we have
$\sqrt{\Theta(\alpha)}\geq\bar{\Lambda}(\alpha-\delta).$
That is,
@@
Taking $n\to\infty$ yields that for any $\delta>0$,
$\frac{\bar{\Lambda}_{k}(\alpha-\delta)}{\sqrt{k}}\leq\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}.$
Applying (5.9) to $(\mathcal{M}^{k},\nu^{\otimes k})$yields that for all $k$,
@@
This is because, on one hand, by the tensorization property of nonlinear log-Sobolev inequality, for all $k$ and $\beta\geq 0$,
(5.12) $\frac{1}{k}\Theta_{k}(k\beta)\geq\breve{\Theta}(\beta);$
on the other hand, by setting the function $f$ in the $k$-dimensional nonlinear log-Sobolev inequality to be the product form $f(x)=\prod_{i=1}^{j}f_{1}(x_{i})\prod_{i=j+1}^{k}f_{2}(x_{i})$ for some functions $f_{1}$ and $f_{2}$ defined on $\mathcal{M}$, the lower bound in (5.12) can be asymptotically approached.
@@
Combining this theorem with Theorem 15 yields that
$\sqrt{\breve{\Theta}(\alpha)}\leq\underline{\Lambda}(\alpha)\leq\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
That is,
$\underline{\Lambda}(\alpha)=\bar{\Lambda}(\alpha)=\sqrt{\breve{\Theta}(\alpha)},$
completing the proof of our main result, Theorem 1.
6. Proof of Theorem 2
@@ 6.1. Inequality (b) in Statement 1
where $r(a)\sim\sqrt{\frac{2}{K_{C}^{+}}\ln\frac{1}{a}}$ as $a\to 0$. Taking limits as $n\to\infty$, we have
@@
Denote $\alpha_{k}=\ln\frac{1}{a_{k}}$, which diverges to $\infty$. We then have
$\frac{1}{2}e^{\alpha_{k}}-1$ $\leq\sqrt{2(K_{IS}^{-}+\epsilon)\alpha_{k}}\int_{0}^{\sqrt{\frac{2}{K_{C}^{+}-\epsilon}\alpha_{k}}}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)\alpha_{k}}t-\frac{K_{CD}}{2}t^{2}\right){\text{\rm d}}t$
$=\sqrt{\pi}e^{\frac{K_{IS}^{-}+\epsilon}{\text{$K_{CD}$}}\alpha_{k}}\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}\alpha_{k}}\Big[\text{erf}\left((\sqrt{\frac{K_{CD}}{K_{C}^{+}-\epsilon}}-\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}})\sqrt{\alpha_{k}}\right)$
$\qquad+\text{erf}\left(\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}\alpha_{k}}\right)\Big],$
where $\text{erf}\left(z\right)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\,\mathrm{d}t$} is the error function. Notice that
$\text{erf}(t)=1-\frac{1+o(1)}{\sqrt{\pi}t}e^{-t^{2}},\;t\to\infty,$
and $\text{erf}(t)=-\text{erf}(-t).$ So, we have
$\frac{1}{2}e^{\alpha_{k}}-1$ $\leq\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}}\Big[-\frac{1+o(1)}{\sqrt{\frac{K_{CD}}{K_{C}^{+}-\epsilon}}-\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}}}$
$\qquad\times\exp\left(\Big(\frac{K_{IS}^{-}+\epsilon}{\text{$K_{CD}$}}-(\sqrt{\frac{K_{CD}}{K_{C}^{+}-\epsilon}}-\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}})^{2}\Big)\alpha_{k}\right)$
$\qquad-\frac{1+o(1)}{\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}}}\Big].$
This inequality holds only if
$\frac{K_{IS}^{-}+\epsilon}{\text{$K_{CD}$}}-\left(\sqrt{\frac{K_{CD}}{K_{C}^{+}-\epsilon}}-\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{CD}}}\right)^{2}\geq 1.$
Solving this inequality and letting $\epsilon\to 0$ yields
$K_{IS}^{-}\geq\left(\frac{1+K_{CD}/K_{C}^{+}}{2}\right)^{2}K_{C}^{+}.$
6.2. $K_{LS}^{+}\leq K_{IS}^{+}$ in Inequality (k)
The part $K_{LS}^{+}\leq K_{IS}^{+}$ in Inequality (k) is exactly the following theorem which is a consequence of Theorem 1.
Theorem 18 (Dimension-Free Isoperimetry).
Assume that (1.5) and (1.8) hold. Then, for any $n\geq 1$, it holds that
(6.3) $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq K_{LS}^{+}.$
Proof.

### v3

@@ 6. The Other Direction: From Standard Isoperimetry to Log-Sobolev
Under Assumption 1, for any $\alpha>0$,
@@
We first prove the first inequality in (6.1). Let $A$ be a set approximately attaining $\Gamma_{n}(e^{-n\alpha},n\tau)$ for $\nu=\nu_{X}=\nu_{Y}$ (given in (3.2)) within a factor $1-\epsilon$ where $\epsilon>0$, i.e., $A$ has probability $e^{-n\alpha}$ and $\nu_{n}(A^{\sqrt{n\tau}})\geq(1-\epsilon)\Gamma_{n}(e^{-n\alpha},n\tau)$. Let $r:=\sqrt{n\tau}$. Since $f(t):=\ln\nu_{n}(A^{t})$ is monotone, it is differentiable almost everywhere on $(0,r)$. As a basic property of monotone functions, $f^{\prime}$ is integrable over $[0,r]$ and $\int_{0}^{r}f^{\prime}(t){\text{\rm d}}t\leq f(r)-f(0)$ (see e.g., [60, Corollary 4 on p.113]), which implies that there is some $r_{0}\in(0,r)$ such that $\frac{f(r)-f(0)}{r}\geq f^{\prime}(r_{0}),$i.e.,
@@
$\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}$ $\geq\lim_{n\to\infty}\frac{\ln[\nu_{n}(A^{\sqrt{n\tau}})/\nu_{n}(A)]}{n\sqrt{\tau}}\geq\limsup_{n\to\infty}\frac{I_{n}(a_{n,\tau})}{a_{n,\tau}\sqrt{n}},$
@@
Since $A$ approximately attains $\Gamma_{n}(e^{-n\alpha},n\tau)$ within a factor $1-\epsilon$, by Statement 2 in Theorem 8 again, it holds that $\liminf_{n\to\infty}-\frac{1}{n}\ln a_{n,\tau}\geq\psi(\alpha,\tau).$ Hence, for any $\delta>0$, $\bar{\Lambda}(\psi(\alpha,\tau)-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$
We can assume that $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$, since otherwise, $\xi(\alpha)=\infty$ and nothing is needed to prove. Hence, given any $\delta>0$, it holds that $\psi(\alpha,\tau)>\alpha-\delta.$ for all sufficiently small $\tau$, which, combined with the inequality above, implies $\bar{\Lambda}(\alpha-\delta)\leq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}}.$ Letting $\tau\downarrow 0$ yields that
@@
$D(\pi_{X|W}\|\nu|\pi_{W})-D(\pi_{Y|W}\|\nu|\pi_{W})=\mathbb{E}_{\pi}[\alpha_{W}-\beta_{W}]$
$\qquad\leq\sqrt{\mathbb{E}_{\pi}[\left(\frac{\alpha_{W}-\beta_{W}}{\gamma_{W}}\right)^{2}]\mathbb{E}_{\pi}[\gamma_{W}^{2}]}.$
@@
We now need the HWI inequality for the Riemannian manifold setting [65, Corollary 20.13]. The original version for Euclidean spaces is due to Otto and Villani [54].
Theorem 17 (HWI Inequality [65, Corollary 20.13]).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that Assumption 1 holds. Then, for any two probability measures $\mu_{0}=\rho_{0}\nu,\mu_{1}=\rho_{1}\nu$ in $P_{2}^{\mathrm{ac}}(\mathcal{M})$ such that $D(\mu_{1}\|\nu)<\infty$ and $\rho_{0}$ is Lipschitz, it holds that
@@
This theorem implies $|\partial^{-}D|(\mu\|\nu)\leq\sqrt{I(\mu\|\nu)}.$ In fact, this is an equality; see [6, Theorem 9.3]. We restrict $\pi_{X|W=i},i\in\{0,1\}$ to be in $P_{2}^{\mathrm{ac}}(\mathcal{M})$. Therefore,
$\xi(\alpha)$ $\leq\inf_{\begin{subarray}{c}\pi_{XW}:\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),\forall i,\\ D(\pi_{X|W}\|\nu|\pi_{W})=\alpha\end{subarray}}\sqrt{I(\pi_{X|W}\|\nu|\pi_{W})}=\sqrt{\breve{\Theta}(\alpha)}.$
Combining this with (6.2) yields that $\bar{\Lambda}(\alpha-\delta)\leq\xi(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$ That is, for $\alpha,\delta>0$, $\bar{\Lambda}(\alpha)\leq\xi(\alpha+\delta)\leq\sqrt{\breve{\Theta}(\alpha+\delta)}.$ Since $\breve{\Theta}(t)$ is continuous in $t>0$, letting $\delta\downarrow 0$ yields $\bar{\Lambda}(\alpha)\leq\sqrt{\breve{\Theta}(\alpha)}.$
@@
By an approximation argument, we may assume that the probability measures $\mu$ in the nonlinear log-Sobolev inequality have smooth density $f:={\text{\rm d}}\mu/{\text{\rm d}}\nu$. Obviously, $\int f\,d\nu=1$. Let $f^{\otimes n}$ be the $n$-fold product of $f$ with itself, which satisfies ${\text{\rm d}}\mu_{n}/{\text{\rm d}}\nu_{n}=f^{\otimes n}$ with $\mu_{n}=\mu^{\otimes n}$. Then, the coarea formula (see e.g., [19, Exercise III.12]) states:
@@
(6.5) $\qquad=\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}^{+}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\geq\eta_{n}\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s,$
@@
$\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s=\int_{0}^{\infty}\int 1\{f^{\otimes n}\geq s,\,s\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}{\text{\rm d}}s$
(6.6) $\qquad=\int(f^{\otimes n}-e^{n(D-\delta)})\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\gamma_{1,n}-\gamma_{2,n},$
@@
$\gamma_{1,n}$ $:=\int f^{\otimes n}\cdot 1\{f^{\otimes n}\geq e^{n(D-\delta)}\}{\text{\rm d}}\nu_{n}=\mu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\},$
@@
By the law of large numbers, $\gamma_{1,n}\to 1$ as $n\to\infty$. By the large deviations theory (or the asymptotic equipartition property in information theory),
@@
where $D_{1-t}(\mu\|\nu)=-\frac{1}{t}\ln\int f^{1-t}d\nu,\;f={\text{\rm d}}\mu/{\text{\rm d}}\nu$ is the (1t)(1-t)-Rényi divergence of $\mu$ w.r.t. $\nu$. Therefore, substituting these into (6.6) yields $\int_{e^{n(D-\delta)}}^{\infty}\nu_{n}\{f^{\otimes n}\geq s\}{\text{\rm d}}s\to 1,\textrm{ as }n\to\infty,$ which, combined with (6.5), implies
@@
$\eta_{n}$ $\geq\inf_{\begin{subarray}{c}s\geq e^{n(D-\delta)}:\\ \nu_{n}\{f^{\otimes n}\geq s\}>0\end{subarray}}\frac{I_{n}(\nu_{n}\{f^{\otimes n}\geq s\})}{\nu_{n}\{f^{\otimes n}\geq s\}}=\frac{I_{n}(a_{n})}{a_{n}},$
where $a_{n}:=\nu_{n}\{f^{\otimes n}\geq e^{n(D-\delta)}\}$, and the equality follows since $I_{n}(a)/a$ is non-increasing in $a$. Hence,
@@
(6.8) $\int|\nabla f^{\otimes n}|\,{\text{\rm d}}\nu_{n}\leq\sqrt{\int f^{\otimes n}\,{\text{\rm d}}\nu_{n}\,\cdot\int\frac{|\nabla f^{\otimes n}|^{2}}{f^{\otimes n}}\,{\text{\rm d}}\nu_{n}}=\sqrt{nI(\mu\|\nu)}.$
Combining (6.4), (6.7), and (6.8) yields that $\sqrt{I(\mu\|\nu)}\geq\bar{\Lambda}(D(\mu\|\nu)-\delta),\,\forall\mu.$ Noting that $\bar{\Lambda}$ is non-decreasing, we have $\sqrt{\Theta(\alpha)}\geq\bar{\Lambda}(\alpha-\delta).$ That is,
@@
Taking $n\to\infty$ yields that for any $\delta>0$, $\frac{\bar{\Lambda}_{k}(\alpha-\delta)}{\sqrt{k}}\leq\bar{\Lambda}(\alpha)\leq\frac{\bar{\Lambda}_{k}(\alpha+\delta)}{\sqrt{k}}.$ Applying (6.9) to $(\mathcal{M}^{k},\nu^{\otimes k})$ yields that for all $k$,
@@
This is because, on one hand, by the tensorization property of nonlinear log-Sobolev inequality, for all $k$ and $\beta\geq 0$, $\frac{1}{k}\Theta_{k}(k\beta)\geq\breve{\Theta}(\beta);$ on the other hand, by setting the function $f$ in the $k$-dimensional nonlinear log-Sobolev inequality to be the product form $f(x)=\prod_{i=1}^{j}f_{1}(x_{i})\prod_{i=j+1}^{k}f_{2}(x_{i})$ for some functions $f_{1}$ and $f_{2}$ defined on $\mathcal{M}$, the lower bound $\breve{\Theta}$ can be asymptotically approached.
@@
Combining this theorem with Theorem 15 yields that $\underline{\Lambda}(\alpha)=\bar{\Lambda}(\alpha)=\sqrt{\breve{\Theta}(\alpha)},$ completing the proof of our main result, Theorem 1.
Appendix A: Proof of Theorem 2
@@ Appendix A-1. Inequality (b) in Statement 1
where $r(a)=\sqrt{\frac{2}{K_{C}}\ln\frac{1}{a}}$. Taking limits as $n\to\infty$, we have
@@
Denote $\alpha_{k}=\ln\frac{1}{a_{k}}$, which diverges to $\infty$. Substituting it to the inequality above and performing change of variables, we then have
$\frac{1}{2}e^{\alpha_{k}}-1$ $\leq\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}\int_{0}^{\sqrt{\frac{2}{K_{C}}}}\exp\left(\sqrt{2(K_{IS}^{-}+\epsilon)}\alpha_{k}t-\frac{K_{CD}}{2}\alpha_{k}t^{2}\right){\text{\rm d}}t.$
Note that $K_{C}\geq K_{IS}^{-}\geq K_{CD}$, which implies $K_{C}\left(K_{IS}^{-}+\epsilon\right)>K_{CD}$. By Laplace’s method, under this condition, the last line above is
$\sim\frac{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}}{\sqrt{\left(K_{IS}^{-}+\epsilon\right)K_{C}}-K_{CD}}\exp\left((2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}})\alpha_{k}\right)$
as $k\to\infty$. It is asymptotically larger than $\frac{1}{2}e^{\alpha_{k}}-1$ only if $2\sqrt{\frac{K_{IS}^{-}+\epsilon}{K_{C}}}-\frac{K_{CD}}{K_{C}}\geq 1$, i.e., $K_{IS}^{-}\geq\left(\frac{1+K_{CD}/K_{C}}{2}\right)^{2}K_{C}$ with letting $\epsilon\to 0$.
Appendix A-2. Inequality (l)
Inequality (l) is exactly the following inequality: for any $n\geq 1$,
(6.14) $\liminf_{a\to 0}\frac{I_{n}(a)}{a\sqrt{2\ln\frac{1}{a}}}\geq\sqrt{K_{LS}^{+}}.$